OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. √169− √50− √8

OpenStudy (anonymous):

Try to split the terms in the square roots into components of which you know the square root. Remember: \[\sqrt{ab}=\sqrt{b}\sqrt{a}\]

OpenStudy (anonymous):

We can do the last term as an example: Lets say \[8=2*4\] Then: \[\sqrt{2*4}=\sqrt{4}*\sqrt{2} = 2*\sqrt{2}\]

OpenStudy (anonymous):

so you do that for every number and then put them together?

OpenStudy (anonymous):

hello?

OpenStudy (mayankdevnani):

@peanutbutter817 what is the the value of this:- \[\huge \bf \sqrt{169}\]

OpenStudy (anonymous):

what is that I cant c it its in word from now

OpenStudy (mayankdevnani):

\[\huge \bf \sqrt{169}=??\]

OpenStudy (anonymous):

I cant c it still r u trying to say √169

OpenStudy (mayankdevnani):

what is the square root of 169 ? @peanutbutter817

OpenStudy (anonymous):

ok its 13

OpenStudy (mayankdevnani):

good

OpenStudy (mayankdevnani):

what is the square root of 50? @peanutbutter817

OpenStudy (anonymous):

7.071067811865476

OpenStudy (mayankdevnani):

no. not so exact. its